Situation: You volunteered to be part of the medical mission in your barangay and in-charge


to get weight of the children. The weights of children have an average of kilograms


and a standard deviation of kilograms.


a. How many children weight 35 kg to 55 kg?


b. How many are in the upper 10% of the population?


c. Estimate the range of the number of children weighing


c.1. 40kg to 50kg


c.2. below 55kg


c.3. between 25kg to 58kg

1
Expert's answer
2022-03-05T06:54:49-0500

Let "X=" weight of children: "X\sim N(\mu, \sigma^2)"

Given "\mu=45kg, \sigma=5kg"

a.

"P(35<X<55)=P(X<55)-P(X\le35)"

"=P(Z<\dfrac{55-45}{5})-P(Z\le\dfrac{35-45}{5})"

"=P(Z<2)-P(Z\le-2)"

"\approx0.97725-0.02275=0.9545"



b.


"1-P(Z\le\dfrac{x-45}{5})=0.1"

"\dfrac{x-45}{5}\approx1.28155"

"x\approx51.408"


c.

c.1.


"P(40<X<50)=P(X<50)-P(X\le40)"

"=P(Z<\dfrac{50-45}{5})-P(Z\le\dfrac{40-45}{5})"

"=P(Z<1)-P(Z\le-1)"

"\approx0.84134-0.15866=0.6827"



c.2.


"P(X<55)=P(Z<\dfrac{55-45}{5})"

"=P(Z<2)\approx0.97725"

c.3.


"P(25<X<58)=P(X<58)-P(X\le25)"

"=P(Z<\dfrac{58-45}{5})-P(Z\le\dfrac{25-45}{5})"

"=P(Z<2.6)-P(Z\le-4)"

"\approx0.99534-0.00003=0.9953"